A 1300 kg aircraft going 40 m/s collides with a 3000 kg aircraft that is parked and they stick together after the collision and are going 12.1 m/s after the collision. If they skid for 151 m before stopping, how long (in seconds) did they skid? Hint: Are the aircraft moving at a constant velocity after the collision or do they experience an acceleration?
Answer #1

Using equation of motion. S= 4x (u+V)xt 151 = 5*( 12.1+0)xt > t = 24.96 s Answer